tdSDoorstope, INC.(DD is a monopolist in the doorstop industry its total cost function 1s oiven by the quadratic function of output Cla) 100-5Q+. The inverse demand function for doorstops Pis given by the linear function PIQ)-55-2Q. Nte that the margnal cost C(Q) is not constant. (Also, it happens to be negative for 0 s Q <2
How much consumer surplus CS and producer surplus PSM and total surplus does DD generate by its profit-maximizing plan? (d) Find the profit-maximizing rate of output, QF, if DD acted like a price-taker (i.e., a perfect competitor) (e) Find the profit-maximizing price set by DD, P, as well as its profit, I, if it acted like a price- taker
Answer #1

Answer:

Given

A)

Cost function C(Q)=100-5Q+Q^2

MC=dC/dQ=2Q-5

MC=2Q-5

P=55-2Q

TR=P*Q=55Q-2Q^2

So MR=dTR/dQ=55-4Q

So for profit maximization for monopolistic firm

MR=MC
55-4Q=2Q-5

Q=10

P=55-2*10=$35

Demand Supply Curve 60 50 MC 20-5 CS 40 35 30 20PS 15 RS 10 P-55-20Q 2.5 5 7.510 12.5 15 17.520 22.5 25 27.5 Quantity

From Above graph we find that CS=area of triangle ABE=0.5*(55-35)*10=$100

From Above graph we find that PS=area of trapzoid EBCD=(35-15)*10+0.5*(15-(-5)*10=$300

Since at MC=-5 , Qs=0

So TS=CS+PS=100+300=$400

B) MC=2Q-5

for price taker firm profit maximization

P=MC

55-2Q=2Q-5

Q=15 units

C) for Q=15 units

P=55-2*15=$25

TR=P*Q=15*25=$375

C(Q)=100-5*15+15^2=$250

So Profit= 375-250=$125